OpenStudy (anonymous):

Change variables and evaluate the new integral. \[\int\limits_{}^{}\int\limits_{R}^{}x ^{2}\sqrt{x+2y}dA\] where \[R=\left\{ (x,y): 0 \le x \le 2, \frac{ -x }{ 2 } \le y \le 1-x\right\}\] Use\[x=2u\] and \[y=v-u.\]

OpenStudy (anonymous):

I've already figured out the new limits of integration which are \[0 \le u \le 1\]\[0 \le v \le 1-u\] and the Jacobian which is 2.

OpenStudy (zarkon):

where are you stuck then?

OpenStudy (anonymous):

I guess the integration.

OpenStudy (zarkon):

so you can setup the integral?

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{1-u} 4u ^{2}\sqrt{2v} \times 2 dvdu\]

OpenStudy (zarkon):

ok

OpenStudy (zarkon):

\[8\sqrt{2}\int\limits_{0}^{1}\left[u^2\int\limits_{0}^{1-u} \sqrt{v} dv\right]du\]

OpenStudy (zarkon):

two calc I integrals

OpenStudy (anonymous):

I get to \[\frac{ 16\sqrt{2} }{ 3 }\int\limits_{0}^{1} u ^{2}(1-u)^{\frac{ 3 }{ 2 }} du\] and I get stuck!

OpenStudy (zarkon):

substitution \(w=1-u\)

OpenStudy (anonymous):

Then what does u^2 become?

OpenStudy (zarkon):

\[w=1-u\] \[u=1-w\] \[u^2=(1-w)^2\]

OpenStudy (anonymous):

Well that makes sense! Thanks!

OpenStudy (anonymous):

Ok but I don't know how to evaluate that integrand either??

OpenStudy (zarkon):

\[(1-w)^2w^{3/2}\] \[(1-2w+w^2)w^{3/2}\] \[w^{3/2}-2w^{5/2}+w^{7/2}\]

OpenStudy (anonymous):

Ok that helps!

OpenStudy (anonymous):

So the answer is\[\frac{ 256\sqrt{2} }{ 945 }\] Thank you!

OpenStudy (zarkon):

yep